OpenStudy (anonymous):

This is so urgent I need help !!! If you have ever swum in a pool and your eyes began to sting and turn red, you are aware of the effects on an incorrect pH level. The pH level measures the concentration of hydronium ions and can be modeled by the function p(t) = –log10 t. The variable t represents the amount of hydronium ions, and p(t) gives the resulting pH level. Water at 25 degrees Celsius has a pH of 7. Anything that has a pH lower than 7 is called acidic, while a pH above 7 is basic, or alkaline. Seawater has a pH just over 8, while lemonade has a pH of approximately 3.

OpenStudy (anonymous):

1. Create a graph of the pH function. Locate on your graph where the pH value is 0 and where it is 1. You may need to zoom in on your graph. I need help creating the function

OpenStudy (horsegirl27):

So a function has to start with f(x)

OpenStudy (anonymous):

Yes

OpenStudy (horsegirl27):

do you know what to do next?

OpenStudy (horsegirl27):

If you know the next step that's good and I can help you place the numbers

OpenStudy (anonymous):

Not even a little bit, I am so confused

OpenStudy (horsegirl27):

Ok

OpenStudy (horsegirl27):

How urgent is this? Do you have time to tell me everything you know about functions? Even if you are not sure it's right tell me

OpenStudy (horsegirl27):

Then I know what I have to do to help

OpenStudy (anonymous):

Im sorry I really don't know. I just know that you have to plug in numbers to solve. For this would we use the format of p(t)=-log10t??? but use f(x) instead?

OpenStudy (anonymous):

@horsegirl27

OpenStudy (horsegirl27):

Ok good guess. BUt no

OpenStudy (anonymous):

Okay sorry.

OpenStudy (horsegirl27):

You replace y with f(x)

OpenStudy (horsegirl27):

no don't be sorry, you don't know but u tried

OpenStudy (horsegirl27):

example: y= 11x+3 f(x)=11x+3

OpenStudy (horsegirl27):

See?

OpenStudy (anonymous):

Okay yeah that makes sense

OpenStudy (horsegirl27):

Then you can give x a value, example: f(x)=11x+3 f(4)=11(4)+3 f(4)=44+3

OpenStudy (horsegirl27):

But we don't need to worry about that yet^

OpenStudy (horsegirl27):

U will probably be annoyed at me for saying this, but I want you to try it. Set it up with what numbers you think are right ok? I know I hate this, because I want people to give me the answers, but I want you to try

OpenStudy (horsegirl27):

the number before x is muliplied by x. Try to find the correct numbers by looking in the problem.

OpenStudy (anonymous):

Okay, let's say... f(x)= 25(x)+x like idk what numbers to use to set it up

OpenStudy (horsegirl27):

brb I will tag someone to help @ganeshie8

OpenStudy (anonymous):

thanks

OpenStudy (horsegirl27):

@jigglypuff314 @Dan20071

OpenStudy (horsegirl27):

@goatsgirl17 help please

OpenStudy (anonymous):

I think I got it.. f(x)= -logx/-log10